OpenStudy (anonymous):

Determine the dimensions of the rectangle of largest area that can be inscribed in a semicircle of radius 3. .. Not sure how to do

OpenStudy (anonymous):

got a picture? it would help to use that the equation of a circle with radius 3 is \[x^2+y^2=9\] and so the upper half would be \[y=\sqrt{9-x^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Well there is a picture but not exactly sure how that would help. The rectangle is in a circle with radius 3. Thats about it. need to find the largest the rectangle can be

OpenStudy (anonymous):

yeah like yours. And from the origin to x is 3, then W and l for the height and length

OpenStudy (anonymous):

so area of that rectangle i drew is \[2x\times \sqrt{9-x^2}\] since the base is 2x and the height is \[\sqrt{9-x^2}\] so find the maximum value of the area function \[A(x)=2x\sqrt{9-x^2}\] i.e. take the derivative, find critical points etc

OpenStudy (anonymous):

Yeah I am not sure if I see it. Dont get how we found the Area to be what you got. I thought it would be just (height)*(width) so would be (x)(\[\sqrt{9-x^2}\]

OpenStudy (zarkon):

it is (height)*(width) the width is 2x